5*7+2=2+5*7 show iz it associative property?
Answer by richard1234(7193)

It's actually an example of the commutative property. Because of the order of operations we do 5*7 first. We can treat 5*7 as a number (in other words, 35) and say that (5*7) + 2 = 2 + (5*7) or 35+2 = 2+35.